### v2

# Space-aware Socioeconomic Indicator Inference with Heterogeneous Graphs

Abstract
Regional socioeconomic indicators are critical across various domains, yet their acquisition can be costly. Inferring global socioeconomic indicators from a limited number of regional samples is essential for enhancing management and sustainability in urban areas and human settlements. Current inference methods typically rely on spatial interpolation based on the assumption of spatial continuity, which does not adequately address the complex variations present within regional spaces. In this paper, we present GeoHG, the first space-aware socioeconomic indicator inference method that utilizes a heterogeneous graph-based structure to represent geospace for non-continuous inference. Extensive experiments demonstrate the effectiveness of GeoHG in comparison to existing methods, achieving an $R^{2}$ score exceeding 0.8 under extreme data scarcity with a masked ratio of 95%.
1 Introduction
Our living space, as a complex system, reflects the interplay between natural laws and social dynamics. Socioeconomic indicators are designed to quantify the societal and economic dynamics of the space and contribute significantly to various domains such as the environment, management, sustainability, and urban computing (Zou et al., 2025; Xie et al., 2020). In practice, these indicators are typically acquired through field investigations (e.g., population census) or deployed equipment (e.g., air monitoring station), with limited budgets resulting in a number of ground truth samples that is much lower than the global demand (Costa et al., 2012; Ross and Reeve, 2003). In this context, Socioeconomic Indicator Inference aims to estimate the values of uninvestigated regions based on limited observations in order to provide a more uniform and comprehensive quantification of the overall geospace.
Figure 1: Insight for space-aware no-continuous indicator inference.
In the past decade, most socioeconomic indicator inference methods, such as the kriging (Brus and Heuvelink, 2007) method, have been developed using geometric approaches for spatial interpolation (Mitas and Mitasova, 1999). While feasible, these methods depend on key assumptions of spatial continuity, spatial stationarity, and isotropic spatial autocorrelation, which imply that the events of the considered spatial indicators are distributed uniformly across the region, regardless of geographical directions or barriers (Mitas and Mitasova, 1999; Tan and Xu, 2014). However, the actual distribution of these indicators is non-continuous and influenced by multiple dependencies. For instance, two adjacent regions may have different population densities due to variations in land environment (Ross and Reeve, 2003), while regions along the same river basin may exhibit similar GDP patterns despite being separated by distance (Li et al., 2024; Wang et al., 2019). Consequently, the interpolated values can significantly diverge from the real scenario. Additionally, the available samples are typically non-uniformly distributed in space, resulting in poor or even ineffective performance of traditional interpolation methods in sparse areas (Li and Heap, 2011).
Figure 2: Illustration of geospatial features and dependencies for spare-aware indicator inference.
To address this issue, it is important to integrate artificial intelligence to uniformly represent features and non-Euclidean relationships within geospace using geospatial data. This representation can then enhance global indicator inference in non-continuous and non-uniform scenarios (see Fig. 1). While significant research on region representation exists (Liu et al., 2023; Wang et al., 2020; Hao et al., 2025; Zhang et al., 2024; Mai et al., 2024), most studies assume redundant samples and train models on large datasets to predict individual samples. The design of a method that comprehensively represents geospatial data for indicator inference remains unexplored.
Introducing such methods presents two main challenges. First, the data reliance for geospace representation must be robust. Existing region representation methods may utilize various data sources, such as trajectories, street-view imagery, or road networks, which might not always be available. Although some studies leverage globally available satellite imagery to generate general spatial representations, their specified vision encoders may perform poorly when trained with only a few samples. Second, comprehensively representing environmental and societal dependencies in geospatial data is non-trivial. As shown in Figure 2, space-aware inference should jointly consider spatial geometry, environmental, and societal dependencies, which extend beyond basic regional features and second-order adjacency to include high-order dependencies across regions. For instance, a certain region may be influenced jointly by a group of non-adjacent regions due to shared geographic or socioeconomic functions.
To address these challenges, we present GeoHG, the first space-aware socioeconomic indicator inference method that utilizes a heterogeneous graph-based structure to jointly represent geospace from spatial geometry, environmental, and societal perspectives for non-continuous inference. For robustness, we employ only satellite imagery and points of interest (POI) to represent the environmental and societal features of the space, as these are among the most accessible data sources. Furthermore, to overcome the limitations of traditional vision encoders, we directly perform semantic segmentation on the satellite imagery to distinguish geo-entities such as water bodies, vegetation, and buildings, enabling a more general and concise representation. For comprehensive geospace representation, we introduce a novel heterogeneous graph that integrates spatial geometrical relations with encoded environmental and societal information. This graph structure effectively captures the non-Euclidean local relationships within the space. Additionally, our heterogeneous formulation between region-entity associations accounts for the high-dimensional societal and environmental dependencies across regions. By uniformly representing the space, we can conduct inference on the graph for the entire region with consistent accuracy, regardless of the distribution of available samples. Overall, our contribution can be summarized as follows:
  • Space-aware Inference. We introduce a new space-aware method for socioeconomic indicator inference that incorporates a joint representation of spatial geometry, environmental, and societal relations. To the best of our knowledge, this is the first work that integrates graph-based space representation for socioeconomic indicator inference.
  • Heterogeneous Graph for Joint Space Representation. We develop a heterogeneous graph structure to model geospace using widely accessible data sources, capturing high-order societal and environmental dependencies through heterogeneous connections.
  • Empirical Evidence. Our experiments demonstrate that GeoHG outperforms existing interpolation methods and region representation models across various downstream tasks. It achieves an $R^{2}$ score exceeding 0.8 under extreme data scarcity with a masked ratio of 95% and performs effectively on inference with highly unevenly distributed samples.
Figure 3: Overall framework of the proposed GeoHG.
2 Formulation
An investigated geospace $\mathcal{R}$ typically comprises multiple regions denoted as =$\mathcal{R}=\left\{r_{1},r_{2},\cdots,r_{N}\right\}$. These regions exhibit unique environmental and societal characteristics, often reflected through various data sources within them. In our study, we represent $r_{i}\in\mathcal{R}$ through the following aspects:
  • Spatial Geometry represents the spatial location and geometrical relations of $r_{i}$. The spatial location can be represented by its center coordinate $(lon^{i},lat^{i})$, and the geometrical relations can be captured by computing the distance between each pair of regions $r_{i}$.
  • Satellite Imagery captured through ongoing earth observation by satellites, provides an overview of the environmental features of $\mathcal{R}$. Each satellite image of a region $r_{i}$ can be expressed as $S_{i}\in\mathbb{R}^{H\times W\times L}$, where $H\times W$ denotes the image’s size and $L$ represents the number of color channels.
  • Point of Interest refers to a specific location or geographic point of significance to society, which includes campuses, businesses, or any other notable locations. POI can be represented as $P=(x,y,c)$, where $x$ and $y$ are the geographical coordinates, and $c$ is the societal category.
Problem Statement. Given a geospace $\mathcal{R}$ with corresponding satellite images $\mathbf{S}=\left\{S_{1},S_{2},\cdots,S_{N}\right\}$ and a set of POIs $\mathbf{P}_{i}=\left\{P_{1},P_{2},\cdots,P_{M}\right\}$, our goal is to construct a graph $\mathcal{G}$ that captures the spatial and contextual characteristics of the space. Within the space graph $\mathcal{G}$, our approach aims to generate precise estimates of socioeconomic indicators $\mathcal{Y}$ for the unlabeled regions $\mathcal{R}_{mask}\subseteq\mathcal{R}$ from the available regions $C_{\mathcal{R}_{mask}}\mathcal{R}$. This process can be formulated as:
$\left(\hat{\mathcal{Y}}^{r_{1}},\hat{\mathcal{Y}}^{r_{2}},\cdots,\hat{\mathcal{Y}}^{r_{k}}\right)\leftarrow\mathcal{F}_{\theta}\left(\mathcal{G}|\mathcal{Y}^{r_{k+1}},\mathcal{Y}^{r_{k+2}},\cdots,\mathcal{Y}^{r_{N}}\right),$
where $\{r_{1},r_{2},\cdots,r_{k}\}\in\mathcal{R}_{mask}$ and we use $\mathcal{M}=k/N$ as the masked ratio. $\mathcal{F}_{\theta}$ is the inference model.
3 Methodology
As illustrated in Figure 3, our GeoHG framework comprises four major components:
  • Space Feature Representation. We first extract basic space features from satellite imagery and POI data, considering the proportion of geo-entities, POI category distribution, and the regions’ positions. The proportions of geo-entities are derived via semantic segmentation of satellite imagery and clustering of POI by categories.
  • Mixed-order Dependencies Representation. To model cross-region spatial and socioeconomic dependencies, we construct a heterogeneous graph connection that reflects the mixed-order relationships from spatial, environmental and societal perspectives.
  • Graph Construction for Geospace. We combine the basic features and cross-region dependencies using a comprehensive graph. The basic features are incorporated into the node vectors, while the mixed-order relations are expressed through the edges and their corresponding weights.
  • Graph-based Indicator Inference. We introduce two approaches to inference regional indicators on the graph: self-supervised graph learning for creating region embeddings for light-weight deployment and end-to-end training to directly inference on the graph for peak performance.
3.1 Space Feature Representation
To effectively encode spatial, environmental, and societal signals into feature embeddings, we extract multi-modal geo-contextual data including satellite imagery, POI data, and the location of the regions.
[…]
$E^{i}_{pos}=(lon^{i}-lon^{0},lat^{i}-lat^{0})\Delta*D^{-1}$ (1)
where $\Delta$ is the transform vector decided by the coordinate system. $D$ denotes the scale of the grids, and for $1\textrm{km}\times 1\textrm{km}$ grids, $D=1$. It is noteworthy that the 1 $1\text{ km}^{2}$ scale can be easily downsampled to other scales for improved compatibility.
Environmental Feature Embedding: We utilize satellite imagery to mine the environmental features and expect the satellite imagery encoding to be efficient and concise. To achieve this, we draw inspiration from the European Space Agency (ESA) with ESA WorldCover Dataset (Zanaga et al., 2022), elaborate in Appendix B, and design a segmentation-based process to encode the satellite images instead of the most commonly used visual encoders such as Resnet. Given a satellite image $S_{i}$ of a region of interest $R_{i}$, we conduct semantic segmentation to obtain a series of geo-entities $\{entity_{1},entity_{2},...,entity_{j}\}_{j=1}^{J}$ within that region, such as developed areas, grass, and trees based on ESA Worldcover, leveraging their area proportion as environmental feature embedding as below:
$E_{env}=\{p_{ent_{1}},p_{ent_{2}},...,p_{ent_{j}}\}_{j=1}^{J}$ (2)
where $p_{ent_{j}}=\frac{A_{j}}{A_{R}}$, and $A_{j}$ is the area occupied by $entity_{j}$, $A_{R}$ is the area of region $R_{i}$, $J$ is the number of environmental entity types. The detail discussion for opting for the segmentation-based method, along with extra validation tests, are elaborated in Appendix C.
Societal Feature Embedding: To effectively represent the societal feature of a region, given POI dataset $D_{POI}$ and the region of interest $R_{i}$, we firstly find all points $D^{R_{i}}_{POI}$($poi_{1},poi_{2},...,poi_{j})$ which are located at $R_{i}$. Then, we count the proportion of different $K$ POI categories in the region $R_{i}$ to get $E_{soc}^{\prime}$ as below:
$E_{soc}^{\prime}=\{p_{poi_{1}},p_{poi_{2}},...,p_{poi_{k}}\}_{k=1}^{K},\quad p_{\text{poi}_{k}}=\frac{|C_{k}\cap D_{POI}^{R_{i}}|}{|D_{POI}^{R_{i}}|}$
where $K$ is the number of POI categories, $C_{k}=\{poi_{j}\in D_{POI}\mid\text{category}(poi_{j})=k\}$. However, directly employing the proportion of different POI categories $E_{soc}^{\prime}$ as societal features can still be problematic since it cannot distinguish between areas with high POI density (usually with a higher degree of socialization) and those with relatively low POI density. Typically, a higher POI density region is more significantly influenced by its social characteristics. Thus, we convert $E_{soc}^{\prime}$ into the final societal feature embedding $E_{soc}$ by applying a Social impact factor $f$ as shown below:
$E_{soc}=f\cdot E_{soc}^{\prime},\quad f=log(D^{R_{i}}_{POI}+1)$ (3)
3.2 Mixed-order Dependencies Representation.
Region indicators are also significantly influenced by various cross-region dependencies. Our objective is to jointly model pairwise second-order spatial relationships between regions, as well as higher-order dependencies among groups of regions that exhibit similar environmental or societal characteristics, using a heterogeneous graph structure.
According to hypergraph theories (Liang et al., 2022; Feng et al., 2019), as elaborated in Appendix D.1, we employ a heterogeneous graph structure to represent cross-region relations. As shown in Figure 4, we utilize heterogeneous nodes in the graph structure as transfer nodes to facilitate high-order message passing (equivalent to hypergraph edges). Given a geospace $\mathcal{R}$, we construct a weighted heterogeneous graph $\mathcal{G}$, which contains three types of nodes: region nodes, environmental entity nodes, and societal entity nodes. Specifically, region nodes correspond to the grid cells of $\{r_{1},r_{2},\cdots,r_{N}\}$; environmental entity nodes represent the $J$ distinct geo-entity classes detected from satellite image $S_{i}$; and societal entity nodes comprise the $K$ different POI categories. The model structure can be found in Figure 3. We will elaborate on how second-order and high-order information is represented.
Figure 4: Illustration of high-order relations representation in GeoHG. Given a hypergraph (a) with high-order relations, we reconstruct it into a heterogeneous graph (b) by adding nodes as higher-order connection channels and get the final representation (c) in $\mathcal{G}$.
Second-Order Relation Representation: Second-order relations capture the pairwise spatial adjacency between regional nodes in $\mathcal{G}$. We construct undirected edges $\mathcal{E}_{RNR}$, Region Nearby Region (RNR) between regional nodes whose corresponding grid cells are spatially adjacent in a $3\times 3$ grid, thereby encoding these local second-order dependencies.
$\mathcal{E}_{RNR,mn}=1,\text{if }\mathcal{V}_{m}\text{ and }\mathcal{V}_{n}\text{ are {adjacent} in geospace}$ (4)
High-Order Relation Representation: Multi-view graphs (Geng et al., 2019; Zou et al., 2025) represent the relations between regions through additional edges and are confined to second-order regional connections. To fully capture the higher-order relations among urban regions, we explicitly model these higher-order interactions utilizing heterogeneous environmental entity nodes $\mathcal{V}_{env}$ and societal entity nodes $\mathcal{V}_{soc}$ within $\mathcal{G}$. Specifically, we construct heterogeneous weighted edges $\mathcal{E}_{ELR}$ and $\mathcal{E}_{SLR}$ through $\mathcal{V}_{env}$ and $\mathcal{V}_{soc}$ to associate each regional node with its constituting environmental entities (e.g., water, vegetation) and societal entities (e.g., educational, commercial POIs). For $p_{ent_{it}}\geq\theta_{Env}$ or $f\cdot p_{poi_{ik}}\geq\theta_{Soc}$:
$\mathcal{E}_{ELR,ij}=1\cdot p_{ent_{ij}},\quad\mathcal{E}_{SLR,ik}=1\cdot f\cdot p_{poi_{ik}}$ (5)
where $p_{ent_{ij}}$ represents the area proportion of geo-entity $j$ at region $i$, $p_{poi_{ij}}$ denotes the quantity proportion of POI category $k$ at region $i$ and $f$ is the social impact factor elaborated in Equation 3. $\theta_{Env}$ and $\theta_{Soc}$ serve as hyperparameters designed to optimize the graph structure for conciseness and efficiency. This edge formulation allows encoding higher-order relationships between regions exhibiting similar environmental/societal attributes, even if they lack spatial proximity. By combining the second-order spatial adjacency edges with these higher-order hyperedge associations, $\mathcal{G}$ holistically represents the mixed-order relational patterns within the region.
3.3 Graph Construction for Geospace
After deriving the basic space features and the inter-regional dependencies encoded in the heterogeneous graph $\mathcal{G}$, we use a model-agnostic graph framework to integrate these information for a comprehensive space representation $E_{space}$. Specifically, this phase can be formulated as:
$E_{space}=\text{GNN}(\mathcal{G}[E_{pos},E_{env},E_{soc}])$ (6)
In the graph, each regional node $v$ in $\mathcal{G}$ is represented with an initial node feature corresponding to its regional feature $E_{pos},E_{env},E_{soc}$ elaborated in Sec 3.1. Note that for simplicity, instead of utilizing hypergraph neural networks (Feng et al., 2019; Yang et al., 2022; Xia et al., 2022), we model hyperedges with the form of complete subgraphs in $\mathcal{G}$. Therefore, we apply Heterogeneous Graph Neural Networks (HGNNs) over $\mathcal{G}$ to update the node representations by iteratively aggregating and transforming multi-hop neighborhood information:
$\mathbf{x}_{v}^{(l+1)}$ $=\text{HGNN}^{(l)}\left(\mathbf{x}_{v}^{(l)},\mathcal{N}(v),\mathcal{G}_{i}\right),\mathbf{x}_{v}^{(l)}\in\mathbb{R}^{d}$ (7)
where $\mathcal{N}(v)$ denotes the neighbors (spatial and higher-order) of node $v$ in $\mathcal{G}$. After $L$ layers of updates, the final node embedding $\mathbf{x}_{v}^{(L)}$ integrates cues from the basic features as well as mixed-order inter-region dependencies.
3.4 Graph-based Indicator Inference
To effectively address various practical demands, we present two inference strategies: self-supervised graph learning, which learns fine-grained, generalizable region representations within the space, enabling lightweight deployment and transferability to diverse downstream tasks; and end-to-end training, which allows for direct inference on the graph to achieve optimal performance.
Self-Supervised Graph Learning: Inspired by CLIP (Radford et al., 2021), we pre-train our model using a contrastive learning paradigm where we maximize the similarity between a region $\mathbf{R}_{i}$ and its corresponding positive region groups $\mathbf{C}_{i}$, while minimizing its similarity with other region groups in the batch. Specifically, we sample regional nodes adjacent to $\mathbf{R}_{i}$ as well as nodes exhibiting similar intra-regional features to construct $\mathbf{C}_{i}$. During training, we leverage a GNN model $\mathcal{M}_{pretrain}$ to obtain embeddings for $\mathbf{R}_{i}$ and all nodes in $\mathbf{C}_{i}$. The embeddings of the sampled positive regional nodes are further pooled to obtain the unified positive region representation $\mathbf{e}_{j}$. The process of generating embeddings for $\mathbf{R}_{i}$ and $\mathbf{C}_{i}$ is conducted as follows:
$\mathbf{e}_{i}=\mathcal{M}_{pretrain}(\mathbf{X}_{i}),\quad\mathbf{e}_{j}=\operatorname{Pool}\left(\mathcal{M}_{pretrain}(\mathbf{X}_{j}),{j\in\mathcal{C}_{i}}\right)$ (8)
We further define a similarity function $f_{\mathbf{score}}$ to measure the similarity between the representations $\mathbf{e}_{i},\mathbf{e}_{j}$. $f_{\mathbf{score}}$ can be a simple dot product or a more complex metric and then employ an InfoNCE-based loss (Oord et al., 2018) to conduct contrastive learning:
$\mathcal{L}_{pretrain}=\mathbb{E}[-\log\frac{\exp\left(f_{\mathbf{score}}\left(\mathbf{e}_{i},\mathbf{e}_{j}\right)\right)}{\sum_{\forall i,n}\exp\left(f_{\mathbf{score}}\left(\mathbf{e}_{i},\mathbf{e}_{n}\right)\right)}]$ (9)
where $f_{\mathbf{score}}(\mathbf{e}_{i},\mathbf{e}_{j})$ represents the score of positive pairs while $f_{\mathbf{score}}\left(\mathbf{e}_{i},\mathbf{e}_{n}\right)$ refers to the scores of negative pairs.
After obtaining the pre-trained model $\mathcal{M}_{pretrain}$ with pre-trained region embedding $\mathbf{E}_{pretrain}$ through self-supervised learning, we fine-tune it with a available samples by adding three trainable linear layers. The weights of these additional linear layers are updated during fine-tuning while the weights of $\mathcal{M}_{pretrain}$ and $\mathbf{E}_{pretrain}$ remain fixed:
$\hat{\mathbf{y}}=\text{MLP}(\mathbf{E}_{pretrain})$ (10)
End-to-End Training: For direct inference of socioeconomic indicators on the graph, the indicator inference can be regarded as a node label prediction task. In practice, we optimize all parameters from scratch using the task’s supervised signal. We adopt the HGNN model $\mathcal{M}$ to obtain embeddings for a given region $i$, which are then fed into a three-linear layer regression head denoted as $\mathbf{Regressor}$. All parameters are updated by minimizing the Mean Square Error (MSE) loss on the available sample nodes. The overall process can be formulated as follows:
$\hat{y}_{i}=\mathbf{Regressor}(\mathcal{M}(\mathbf{X}_{i})),\quad\mathcal{L}_{\text{MSE}}=\frac{1}{N}\sum_{i=1}^{N}(\hat{y}_{i}-y_{i})^{2}$
4 Experiments
4.1 Experimental Setup
Datasets. The datasets used in this paper include satellite imagery, POI information and five region indicators (Population, GDP, Night Light, Carbon, $PM_{2.5}$) located in four representative cities in China: Beijing, Shanghai, Guangzhou, and Shenzhen. All indicators are provided at 1 $km^{2}$ scale by data servers. The satellite images and POI data are collected from Google Map API according to geospatial grids. The entity segmentation results for satellite imagery are collected from the European Space Agency. The overall dataset statistics can be listed in Table 1 and more information about region indicators can be found in Appendix E. We randomly mask 75% data as unlabeled samples for test, the available samples are divided to 80% for train and 20% for validation.
Table 1: Dataset statistics.
Dataset Coverage in Geospace Satellite Image POI Data
Bottom-left Top-right
Beijing 39.75°N, 116.03°E 40.15°N, 116.79°E 4,277 709,232
Shanghai 30.98°N, 121.10°E 31.51°N, 121.80°E 5,292 808,957
Guangzhou 22.94°N, 113.10°E 23.40°N, 113.68°E 8,540 805,997
Shenzhen 22.45°N, 113.75°E 22.84°N, 114.62°E 5,150 717,461
[…]
Methods GeoHG GeoHG-SSL UrbanVLP GeoStructural PG-SimCLR UrbanCLIP IDW UK
Metric $\text{R}^{2}$ MAE RMSE $\text{R}^{2}$ MAE RMSE $\text{R}^{2}$ MAE RMSE $\text{R}^{2}$ MAE RMSE $\text{R}^{2}$ MAE RMSE $\text{R}^{2}$ MAE RMSE $\text{R}^{2}$ MAE RMSE $\text{R}^{2}$ MAE RMSE
Beijing Carbon 0.894 0.112 0.211 0.887 0.161 0.224 0.717 0.363 0.467 0.705 0.398 0.482 0.402 0.651 0.764 0.604 0.558 0.598 0.734 0.367 0.489 0.745 0.355 0.454
Population 0.834 0.271 0.351 0.830 0.282 0.374 0.695 0.404 0.513 0.701 0.402 0.504 0.451 0.964 1.117 0.431 0.552 0.669 0.689 0.385 0.517 0.701 0.381 0.499
GDP 0.617 0.336 0.567 0.614 0.331 0.581 0.556 0.416 0.650 0.567 0.401 0.609 0.247 0.768 1.254 0.315 0.539 0.864 0.555 0.414 0.619 0.411 0.621 0.409
Night Light 0.861 0.239 0.311 0.850 0.244 0.353 0.531 0.394 0.629 0.488 0.429 0.657 0.369 0.404 0.728 0.420 0.457 0.700 0.668 0.304 0.414 0.681 0.301 0.384
$PM_{2.5}$ 0.921 0.064 0.160 0.920 0.065 0.169 0.641 0.484 0.954 0.694 0.306 0.482 0.398 0.624 0.845 0.533 0.556 0.699 0.791 0.139 0.271 0.809 0.129 0.234
Shanghai Carbon 0.915 0.157 0.290 0.912 0.162 0.312 0.716 0.392 0.529 0.688 0.413 0.557 0.298 0.712 0.914 0.671 0.426 0.569 0.696 0.395 0.563 0.713 0.388 0.532
Population 0.936 0.161 0.244 0.928 0.172 0.251 0.593 0.471 0.607 0.613 0.456 0.583 0.315 0.731 0.959 0.456 0.557 0.748 0.617 0.457 0.607 0.686 0.342 0.498
GDP 0.778 0.323 0.468 0.767 0.331 0.477 0.330 0.595 0.816 0.377 0.553 0.721 0.294 0.767 1.052 0.326 0.587 0.807 0.319 0.611 0.837 0.331 0.544 0.811
Night Light 0.898 0.222 0.311 0.891 0.234 0.347 0.457 0.494 0.667 0.442 0.517 0.685 0.308 0.566 0.768 0.387 0.511 0.709 0.444 0.571 0.756 0.464 0.517 0.729
$PM_{2.5}$ 0.866 0.120 0.379 0.836 0.150 0.394 0.486 0.497 0.654 0.527 0.398 0.592 0.303 0.617 0.895 0.444 0.518 0.774 0.592 0.342 0.554 0.643 0.291 0.542
Guangzhou Carbon 0.885 0.219 0.336 0.884 0.209 0.371 0.698 0.385 0.514 0.681 0.497 0.529 0.422 0.708 0.708 0.585 0.444 0.603 0.675 0.415 0.503 0.754 0.370 0.433
Population 0.871 0.244 0.274 0.855 0.255 0.299 0.665 0.441 0.481 0.687 0.433 0.485 0.303 0.954 0.972 0.533 0.567 0.687 0.674 0.471 0.483 0.695 0.433 0.438
GDP 0.715 0.371 0.532 0.712 0.366 0.574 0.436 0.541 0.764 0.439 0.533 0.699 0.282 0.897 1.264 0.440 0.546 0.762 0.411 0.625 0.748 0.476 0.611 0.763
Night Light 0.871 0.234 0.378 0.854 0.248 0.391 0.577 0.418 0.573 0.574 0.415 0.581 0.435 0.433 0.627 0.483 0.478 0.633 0.442 0.451 0.604 0.476 0.402 0.598
$PM_{2.5}$ 0.833 0.133 0.403 0.822 0.158 0.414 0.638 0.462 0.624 0.652 0.461 0.597 0.315 0.542 0.741 0.56 0.514 0.694 0.671 0.394 0.559 0.696 0.380 0.503
Shenzhen Carbon 0.926 0.128 0.290 0.912 0.162 0.304 0.659 0.418 0.568 0.647 0.431 0.581 0.257 0.683 0.816 0.562 0.483 0.571 0.241 0.577 0.726 0.189 0.634 0.887
Population 0.892 0.165 0.244 0.879 0.173 0.261 0.790 0.343 0.448 0.797 0.314 0.390 0.311 0.758 0.892 0.527 0.592 0.610 0.719 0.354 0.455 0.775 0.342 0.421
GDP 0.798 0.297 0.468 0.767 0.331 0.489 0.532 0.448 0.676 0.517 0.455 0.699 0.307 0.895 1.003 0.508 0.464 0.693 0.534 0.417 0.625 0.553 0.404 0.623
Night Light 0.942 0.149 0.245 0.939 0.155 0.268 0.457 0.459 0.667 0.445 0.488 0.682 0.454 0.358 0.588 0.387 0.511 0.709 0.413 0.513 0.719 0.396 0.528 0.758
$PM_{2.5}$ 0.906 0.116 0.298 0.905 0.117 0.331 0.566 0.494 0.598 0.597 0.451 0.491 0.323 0.613 0.883 0.430 0.586 0.762 0.653 0.398 0.426 0.749 0.345 0.391
Baselines. For comparison, we select two traditional interpolation methods (Inverse distance weighted (IDW) (Zimmerman et al., 1999; Lu and Wong, 2008) and Universal Kriging (UK) (Brus and Heuvelink, 2007)), the state-of-the-art LLM-enhanced methods ( UrbanCLIP (Yan et al., 2024)) and conventional socioeconomic indicator prediction methods (UrbanVLP (Hao et al., 2024), PG-SimCLR (Xi et al., 2022)) and graph-based approach (GeoStructual (Li et al., 2022)). Details about baselines is provided in Appendix F.
[…]
4.2 Comparison with State-of-the-Art Methods
We conduct an empirical assessment of various models across the four datasets, and the results of each model are detailed in Table 2. From these tables, we have two key findings: 1)Both GeoHG and GeoHG-SSL outperform all competing baselines over the 5 datasets for 4 cities. For instance, GeoHG surpasses the previous SOTA performance in Beijing, achieving $R_{2}$ improvements of +16.7%, +14.9%, +6.1%, +37%, +33% in Carbon, Population, GDP, Night Light and $PM_{2.5}$ tasks respectively. Consistent trends are observed across other cities, underscoring GeoHG’s stable accuracy. 2) Most socioeconomic indicator prediction methods perform even worse than traditional spatial interpolation methods on indicator inference tasks. GeoHG captures both spatial geometry and the underlying socioeconomic dynamics of geospace, resulting in lower data reliance compared to existing methods.
Figure 5: Visualization of Population Inference Results.
Besides, to better demonstrate the advantages of GeoHG’s non-continuous inference on large geospace, we evaluate the model against baseline methods on population inference in an expanded area of Shenzhen, using only 5% available data (863 points) to infer the population of remaining 16,401 $km^{2}$ space, as illustrated in Figure 5. The model accurately captures the true distribution, substantially outperforming traditional spatial interpolation methods like IDW and UK, which deviate significantly. This highlights that traditional methods, which rely solely on spatial relationships for interpolation, fail to model the intricate socio-environmental dependencies that are critical for characterizing the distribution of real-world socioeconomic indicators. In contrast, the effective joint modeling of geospace by GeoHG enables accurate indicator inference more than only spatial geometric dimensions.
Figure 6: Results of Ablation Study on $R^{2}$ Metric.
4.3 Ablation Study & Interpretation Analysis
Effects of Core Components. To examine the effectiveness of each core module, we conducted an ablation study based on the following variants: a) w/o Env, which excludes environmental information from the satellite imagery for embedding; b) w/o Soc, which omits POI entities for embedding; and c) w/o Pos, which does not utilize the corresponding location information of the regions. The $R^{2}$ results for five tasks in four cities are shown in Figure 6. We observe that removing location information significantly degrades performance in all tasks. Additionally, environmental features have a substantial impact on model accuracy for the PM2.5, Night Light, and GDP tasks. These results align with the intrinsic distribution characteristics of the task datasets. For instance, PM2.5 and night light data are primarily related to environmental elements such as vegetation and infrastructure. Conversely, data for carbon estimation is mainly derived from societal activity metrics. The analysis of the four cities reveals distinct distribution patterns across the five key indicators, highlighting the heterogeneity in geospatial characteristics. However, the robust performance of our model, which is constructed through the synthesis of three distinct perspectives, remains evident.
Effect of Graph Construction. Our framework utilizes heterogeneous graph and hyperedges between regions to reflect the geospatial high-order relations. To validate the effectiveness of high-order relation representation, we compare GeoHG against two variants: GeoHG-MLP which discards graph structure and relies only on intra-region feature representation, employing MLP for regression; GeoHG-Mono which keeps edges in adjacent regions while discarding hyperedges for high-order relations. The results presented in Figure 6 indicated that our GeoHG significantly benefits from efficient representation of complex mixed-order dependiencies in geospace, thereby resulting in enhanced performance.
Figure 7: Visualization of Geospatial Correlations learned by GeoHG. $(x,y)$ represents the numbering in the latitude and longitude.
Interpretation of Mixed-order Geospatial Correlations. To investigate the power of heterogeneous graph structure in capturing mixed-order dependencies for inference, inspired by GNNExplainer (Ying et al., 2019), we depict the trend of learned geospatial dependency for Carbon Emission inference in Guangzhou. By selecting the top $N$ important nodes by their weight in the GNN prediction process, as illustrated in Figure 7, we observe that GeoHG effectively captures the high-order correlations between the target region and remote region groups through message passing within the heterogeneous graph structure. The details about GNNExplainer and experiment results are introduced in Appendix F.2.
4.4 Qualitative Demonstration
To qualitatively assess the effectiveness of GeoHG in modeling geospace, we illustrate the regional dependency distribution by computing the cosine similarity obtained in GeoHG-SSL between different regions. As shown in Figure 8, we compare the relevant real-world situations of regions to examine our method’s ability to capture features and dependencies within the space. Due to page limitation, we introduce thr detailed explanation of the visualization results in Appendix G.
Figure 8: Regions’ similarity in GeoHG of Guangzhou. Darker indicate higher similarity with the region of interest.
4.5 Robustness Testing on Data Efficiency
To validate the robustness of GeoHG for inference with extremely limited data, we evaluate the model’s performance under masked ratios of 1%, 5%, 10%, and 15%, with results shown in Table 3. The results demonstrate the model’s strong data efficiency; with only 5% of the data, the model outperforms previous state-of-the-art socioeconomic indicator prediction models, such as UrbanVLP (Hao et al., 2024) and GeoStructural (Li et al., 2022), even though these models were trained on the entire training dataset across all tasks as described in their original papers.
Table 3: Data Efficiency of GeoHG in indicator inference. We take the average of $R^{2}$ results in 4 cities.
Masked Ratio Carbon Population GDP Light PM2.5
99% 0.720 0.713 0.535 0.703 0.694
95% 0.835 0.818 0.623 0.844 0.832
90% 0.850 0.846 0.635 0.864 0.856
80% 0.877 0.861 0.724 0.876 0.867
5 Related Work
Deep Learning for Geospace Embedding. Numerous prior studies (Zhang et al., 2021; Zhou et al., 2023; Huang et al., 2023; Wang et al., 2020; Li et al., 2022) have tackled spatial embedding in data-driven solutions. For example, Zhang et al. (2021) used a multi-view graph representation that integrates POI data and mobility patterns, while Zhou et al. (2023) applied a prompt learning technique combining both POI and mobility data. Recent models like PG-SimCLR (Xi et al., 2022) and UrbanCLIP (Yan et al., 2024) have succeeded in profiling urban regions through satellite images. However, these models still struggle with indicator inference from limited samples and do not sufficiently explore the joint representation of geospace from geometric, environmental, and societal perspectives.
Socioeconomic Indicator Inference. Socioeconomic indicator inference is based on spatial interpolation methods, including Inverse Distance Weighting (IDW) (Zimmerman et al., 1999; Lu and Wong, 2008) and Kriging (Brus and Heuvelink, 2007), which are commonly used for this purpose. IDW is a straightforward method that weights sample points inversely by distance, but it does not consider spatial correlation and assumes a monotonous decrease in influence with distance (Mitas and Mitasova, 1999). In contrast, Kriging is a more complex method that incorporates both distance and spatial correlation, utilizing variograms and covariance functions to minimize prediction errors and provide uncertainty estimates. However, Kriging is computationally intensive for large spatial domains and relies on assumptions of spatial stationarity and isotropy (Leal-Villaseca et al., 2025). These widely-used interpolation methods are built on the assumptions of spatial continuity and isotropy.
6 Conclusion and Future Work
In this paper, we introduce GeoHG, a space-aware method for socioeconomic indicator inference that employs a heterogeneous graph-based geospace representation framework. Our approach enhances indicator inference from environmental and societal perspectives, as well as higher-order cross-region correlations. Extensive experiments demonstrate that GeoHG outperforms existing methods while maintaining strong performance even with limited data. Its minimal data reliance and double deployment strategies facilitate easy implementation for various domains. In the future, we plan to integrate advanced graph geometric techniques to accommodate more complex spatial structures within our graph framework, such as spherical space representation. This enhancement will allow the method to handle larger or global-scale inference where planar geometry is inadequate. Additionally, we aim to refine our code into a plug-and-play toolkit for indicator inference and interpolation, benefiting researchers in various fields.
Supplementary for:
“Space-aware Socioeconomic Indicator Inference with Heterogeneous Graphs”
Appendix A More Related Work
Multimodal Learning in GeoAI. Multimodal learning involves integrating data from various modalities to enhance model performance in various tasks [Baltrušaitis et al., 2018; Guo et al., 2019; Gao et al., 2020]. Traditional geospatial embedding learning mostly relies on task-specific supervised methods [Yuan et al., 2012; Zheng, 2015], which are limited by their dependence on domain expertise and inability to adapt to new tasks [Zou et al., 2025; Guo et al., 2021]. For unified and foundation region embedding, recent studies [Jenkins et al., 2019; Zhang et al., 2021] focus on learning general geospatial embeddings from a multimodal perspective. Nonetheless, efficiently integrating multimodal data from intricate geospaces, while maintaining generalizability, poses a significant challenge [Shang et al., 2020b; Zou et al., 2025; Rao et al., 2024; Shang et al., 2020a]. Specifically, conventional vision encoders (including multimodal large language models) trained on natural images find it challenging to grasp the geo-semantic details of satellite images, while the social impact from POIs in geospace is often overlooked. Additionally, popular modality fusion approaches in multimodal learning, like CLIPs [Hao et al., 2024; Yan et al., 2024], fall short in comprehending the cross-modal interactions between geo-entities within geospace. In this research, drawing inspiration from the success of semantic segmentation of satellite imagery [Zanaga et al., 2021] and heterogeneous graph methodologies, we propose an integrated geospatial embedding learning framework based on HG architecture that merges visual entities and POI attributes of geospace while considering mixed-order geospatial relationships with entities and regions.
Appendix B More Introduction of ESA WorldCover Dataset
Figure 9: Visualization of ESA WorldCover Dataset.
The ESA WorldCover dataset  [^1] https://esa-worldcover.org/en represents a groundbreaking advancement in land use/land cover mapping, offering freely accessible, high-resolution (10 m) global coverage based on satellite imagery. Inspired by the 2017 WorldCover conference  [^2] https://worldcover2017.esa.int/, the European Space Agency (ESA) launched the WorldCover project and the primary accomplishment of this endeavor was the introduction in October 2021 of a freely accessible global WorldCover Dataset at a groundbreaking 10 m resolution for the year 2020 [Zanaga et al., 2021, 2022]. This dataset leverages satellite imagery from both Sentinel-1 and Sentinel-2, and encompasses 11 distinct geo-entity categories, shown in Figure 9. It has also undergone rigorous independent validation by Wageningen University (for statistical accuracy) and the International Institute for Applied Systems Analysis (IIASA) (for spatial accuracy), attaining a notable global overall accuracy of approximately 75% [Venter et al., 2022].
Figure 10: Examples of Utilized Data in Our Paper.
Figure 11: Geo-entity Segmentation for Satellite Images based on ESA WorldCover.
WorldCover dataset is continuously updated and revised by ESA with a highly efficient processing pipeline and scalable infrastructure. The core model is trained on a total of 2,160,210 Sentinel-2 images and is able to process the whole world in less than 5 days [Zanaga et al., 2021, 2022]. The revised version we utilized in this paper was released on 28 October 2022, which elevated the global overall accuracy to 76.7% [Zanaga et al., 2022]. This version is free of charge to the entire community and widely accepted by the United Nations Convention to Combat Desertification (UNCCD), the World Resources Institute (WRI), the Centre for International Forestry Research (CIFOR), the Food and Agriculture Organization (FAO) and the Organisation for Economic Co-operation and Development (OECD). We visualized 3 examples of the WorldCover data unitized in this paper in Figure 10.
Appendix C Entity Segmentation for Satellite Image Encoding
C.1 Semantic segmentation-based approach
It is apparent that traditional vision encoders are designed and trained on natural images, which are considerably different from satellite images. Understanding the complex and highly specialized geo-semantic information in satellite imagery is a challenging task.
Luckily, satellite images display a significant level of structured organization concerning semantic content, in contrast to natural images, which contain a vast array of diverse information. To accurately analyze a satellite image, one needs to focus solely on the geological entities it includes, the extent of these entities, and their positions within the geospace. Therefore, using an entity segmentation method to encode satellite images is very promising, as it can directly offer crucial information regarding the entities and their spatial distribution.
In our proposed method, GeoHG, we employ an entity segmentation-based framework as the backbone for encoding satellite imagery, as opposed to the presently prevalent vision encoders, such as CNN and Transformer. To implement this approach, we utilize the ESA WorldCover Dataset, as discussed in Appendix B, due to its extensive validation by geographical researchers and demonstrated robust global generalization capabilities. As depicted in Figure 11, we directly retrieve the geo-entity segmentation results from the WorldCover Dataset for the input satellite imagery. Subsequently, we determine the proportion of geo-entities within the relevant region and construct the environmental feature embedding $E_{env}$.
An additional advantage of our proposed approach lies in its ability to explicitly construct high-order connections among regions, geo-entities, and other regions based on the semantic content of satellite imagery. This is not feasible with traditional encoders. Moreover, since all segmentation results are readily available, this method significantly reduces computational overhead. We have observed that the segmentation-based approach dramatically decreases our training time. For example, the training time for one epoch on the Population dataset in Beijing was reduced substantially from $\approx 25minutes$to $\approx 0.2seconds$.
Table 4: Validation Experiment Results. The bold/underlined font means the best/the second-best result.
Dataset Beijing
Carbon Population GDP Night Light $PM_{2.5}$
Model $R^{2}$ MAE RMSE $R^{2}$ MAE RMSE $R^{2}$ MAE RMSE $R^{2}$ MAE RMSE $R^{2}$ MAE RMSE
Autoencoder 0.298 0.565 0.844 0.168 0.667 0.918 0.171 0.822 1.251 0.276 0.643 0.784 0.209 0.659 0.840
ResNet-18 0.394 0.577 0.805 0.266 0.623 0.857 0.285 0.699 1.024 0.348 0.576 0.734 0.341 0.589 0.821
PG-SimCLR 0.442 0.631 0.754 0.471 0.964 1.117 0.277 0.768 1.254 0.369 0.404 0.728 0.398 0.624 0.845
UrbanCLIP 0.664 0.528 0.598 0.461 0.552 0.669 0.355 0.539 0.864 0.420 0.457 0.700 0.533 0.556 0.699
GeoSegment 0.504 0.513 0.745 0.562 0.514 0.651 0.531 0.445 0.675 0.606 0.470 0.629 0.468 0.535 0.761
Dataset Shanghai
Carbon Population GDP Night Light $PM_{2.5}$
Model $R^{2}$ MAE RMSE $R^{2}$ MAE RMSE $R^{2}$ MAE RMSE $R^{2}$ MAE RMSE $R^{2}$ MAE RMSE
Autoencoder 0.230 0.532 0.771 0.166 0.742 0.898 0.197 0.844 1.511 0.164 0.617 0.729 0.243 0.691 0.942
ResNet-18 0.326 0.465 0.763 0.279 0.627 0.797 0.289 0.711 1.137 0.244 0.571 0.756 0.292 0.642 0.954
PG-SimCLR 0.298 0.712 0.914 0.315 0.731 0.959 0.294 0.767 1.052 0.308 0.566 0.768 0.303 0.617 0.895
UrbanCLIP 0.671 0.426 0.569 0.461 0.552 0.748 0.326 0.587 0.807 0.387 0.511 0.709 0.444 0.518 0.774
GeoSegment 0.791 0.242 0.452 0.812 0.264 0.396 0.578 0.473 0.650 0.684 0.415 0.559 0.656 0.356 0.582
Dataset Guangzhou
Carbon Population GDP Night Light $PM_{2.5}$
Model $R^{2}$ MAE RMSE $R^{2}$ MAE RMSE $R^{2}$ MAE RMSE $R^{2}$ MAE RMSE $R^{2}$ MAE RMSE
Autoencoder 0.254 0.570 0.733 0.195 0.753 0.928 0.176 0.811 1.463 0.176 0.602 0.798 0.196 0.780 0.903
ResNet-18 0.375 0.515 0.673 0.274 0.671 0.831 0.251 0.725 1.048 0.242 0.551 0.704 0.231 0.694 0.859
PG-SimCLR 0.422 0.708 0.708 0.303 0.954 0.972 0.282 0.897 1.264 0.435 0.433 0.627 0.315 0.542 0.741
UrbanCLIP 0.585 0.444 0.603 0.533 0.567 0.687 0.440 0.546 0.762 0.483 0.478 0.633 0.560 0.514 0.694
GeoSegment 0.490 0.519 0.657 0.685 0.409 0.563 0.536 0.488 0.681 0.618 0.453 0.620 0.493 0.435 0.711
Dataset Shenzhen
Carbon Population GDP Night Light $PM_{2.5}$
Model $R^{2}$ MAE RMSE $R^{2}$ MAE RMSE $R^{2}$ MAE RMSE $R^{2}$ MAE RMSE $R^{2}$ MAE RMSE
Autoencoder 0.189 0.634 0.887 0.175 0.772 0.921 0.119 0.884 1.733 0.166 0.608 0.758 0.149 0.645 0.891
ResNet-18 0.241 0.577 0.726 0.299 0.654 0.855 0.234 0.817 1.125 0.243 0.543 0.719 0.273 0.598 0.826
PG-SimCLR 0.257 0.683 0.816 0.311 0.758 0.892 0.307 0.895 1.003 0.454 0.488 0.682 0.597 0.451 0.491
UrbanCLIP 0.562 0.483 0.571 0.527 0.592 0.610 0.508 0.464 0.693 0.387 0.511 0.709 0.460 0.586 0.762
GeoSegment 0.542 0.518 0.698 0.602 0.438 0.602 0.658 0.411 0.587 0.738 0.386 0.512 0.471 0.480 0.752
C.2 Validation experiment
To validate the effectiveness of our segmentation-based approach. We devise a validation experiment on predicting five indicators for 4 cities by taking the segmentation-based geo-entity proportion $E_{env}$ as the only input (named GeoSegment for convenience in Tabel 4). In comparison to the state-of-the-art LLM-enhanced satellite imagery encoder UrbanCLIP [Yan et al., 2024], POI information-enhanced imagery encoder PG-SimCLR [Xi et al., 2022] and two traditional image encoders ResNet-18 and AutoEncoder. These baselines are detailed and introduced in Appendix F.
From Table 4, it is evident that our semantic segmentation-based approach overwhelmingly outperforms all baseline models in almost every task. This includes the state-of-the-art LLM-enhanced satellite imagery encoder, UrbanCLIP [Yan et al., 2024], and the POI information-enhanced encoder, PG-SimCLR [Xi et al., 2022]. Notably, GeoSegment achieves these significant advantages without relying on any additional information, such as LLM descriptions or POI data. This demonstrates the superiority of our method in encoding satellite images.
Appendix D Theoretical Analysis and Details of GeoHG
D.1 Graph theory and our motivation
D.1.1 Basic graph theory.
Graphs provide a natural abstraction to represent structured relationships between entities (nodes) and their attributes (features). Within this framework, knowledge and information are organized via connectivity patterns encoded by edges. A key advantage of graph-structured representations is that nodes can reason about their representations not just from their attributes, but also by recursively aggregating information from their neighbors. This allows graphs to effectively capture complex relational dependencies and contextual patterns.
Graph Neural Networks (GNNs) have emerged as a powerful paradigm for learning representations on graph-structured data by leveraging the message-passing mechanism. The core idea of this mechanism involves each node recursively aggregating representation vectors from its local neighborhood, allowing it to accumulate information from an expanding neighborhood scope across iterations [Wu et al., 2020]. Formally, we can represent this process through the message-passing equation:
$m_{u}$ $=\operatorname{Aggregate}\left(f_{v},v\in\mathcal{N}_{u}\right),$ (11)
$f_{u}^{\prime}$ $=\operatorname{Update}\left(m_{u},f_{u}\right),$
here, $f_{u}$ represents the original representation of node $u$ while $f_{u}^{\prime}$ denotes its new representation after one iteration, $\mathcal{N}_{u}$ is its set of neighboring nodes. Through this iterative process, GNNs can effectively capture dependencies spanning the entire graph topology, enabling them to learn highly expressive representations that fuse localized features with broader structural context.
D.1.2 Hypergraph theory.
Traditional graph models can only connect pairs of vertices within edges, thereby facing limitations in representing higher-order multi-way relationships beyond simple pairwise associations. In real-world scenarios, relationships between data typically go beyond simple pairwise links and involve complex multi-element patterns. For example, in a transportation network, a route often spans multiple cities or locations rather than just directly connecting two endpoints, which is difficult to capture using traditional graphs. To overcome this, hypergraph theory generalizes the graph formulation by introducing hyperedges that can connect any number of vertices, naturally facilitating the representation of higher-order multi-way relationships, defined as:
$G=(V,E);V=\{v_{1},v_{2},...,v_{N}\},E=\{\mathcal{H}_{1},\mathcal{H}_{2},...,\mathcal{H}_{M}\}$ (12)
where $G$ represents a hypergraph and V is the set of nodes. $E=\{\mathcal{H}_{1},\mathcal{H}_{2},...,\mathcal{H}_{M}\}$ is the set of hyperedges representing connectivity among nodes and hyperedge $\mathcal{H}_{m}$ is a subset of $V$. Revisiting the transportation route example, we can model each city/location as a node and the route passing through multiple locations as a hyperedge connecting all the corresponding nodes, precisely encoding the dependencies between the route and the groups of locations it traverses.
D.2 Details of GeoHG structure
Table 5: Major entities in GeoHG
Entity Num Examples
Region N $1\textrm{km}\times 1\textrm{km}$ grid cells from the geospace. The number of region entities depends on the number of grids within a given geographical range.
Environment 9 Tree cover, Shrubland, Grassland, Cropland, Built-up, Bare/sparse vegetation, Permanent water bodies, Herbaceous wetland, Moss and lichen
Society 14 Food and Beverage, Transportation facilities, Shopping spend, Science, education and culture, Companies, Recreation, Financial Institutions, Tourist Attractions, Life services, Car related, Sports and fitness, Hotel accommodation, Healthcare, Commercial residence
Table 6: Major relations in GeoHG
Relation Weight Head & Tail Entity Abbrev.
Region Nearby Region - (Region, Region) RNR
Environmental Entity Locates at Region the area occupied by entity (Environmental Entity, Region) ELR
Societal Entity Locates at Region the transformed proportion of POI category (Societal Entity, Region) SLR
We introduce a novel heterogeneous hypergraph representation, termed GeoHG, to encode the high-order relational structures within geospatial data. GeoHG comprises three types of nodes: regional nodes, environmental entity nodes, and societal entity nodes, along with the relations between them. Specifically, regional nodes correspond to $1km\times 1km$ grid cells from the geographic space; environmental entity nodes represent the 9 distinct geo-entity classes detected from satellite imagery; societal entity nodes constitute the 14 different point-of-interest (POI) categories within the regions.
We construct hyperedges connecting multiple regional nodes through the environmental and societal entity nodes, enabling the explicit modeling of higher-order inter-region relationships induced by environmental or societal factors. The semantics of the node types and their relationships within GeoHG are illustrated in Tables 5 and 6, respectively.
By introducing hyperedges that can associate arbitrary subsets of regional nodes, GeoHG effectively captures the complex high-order dependencies spanning environmental conditions, geographic entities, social dynamics, and their intricate interplay across different spatial regions. This enriched relational structure encodes valuable contextual signals that are integrated into the region representations learned by GeoHG, ultimately benefiting a wide range of downstream geospatial analytics tasks.
Appendix E Dataset and Experiment Settings
E.1 Dataset details.
We employ five representative tasks located in four cities in China: Beijing, Shanghai, Guangzhou and Shenzhen. Population, GDP, and Night Light tasks reflect anthropogenic activities, whereas Carbon and Temperature tasks characterize the natural environment. We conduct data preprocessing according to [Yan et al., 2024]. The detailed information of datasets is listed below:
  • Carbon Emissions: This dataset incorporates anthropogenic CO2 emission estimates sourced from the Open Data Inventory (ODIAC) [^3] https://odin.opendatawatch.com/ for the year 2022, spatially aligned with our 1 km2 satellite image grids (emissions quantified in tons).
  • Population: This dataset is obtained through WorldPop [^4] https://www.worldpop.org/’s population distribution data for 2020, with counts representing the number of citizens per region.
  • GDP: This dataset includes Gross Domestic Product (GDP) statistics reflecting China’s regional economic development patterns from Zhao et al. [2017].
  • Night Light: As a proxy for human activity intensity, a key driver of urban evolution, we leverage nighttime light imagery data from Zhong et al. [2022] in 2020.
  • $PM_{2.5}$: The $PM_{2.5}$ dataset is sourced from ChinaHighPM2.5 dataset [WEI Jing, 2024]. This dataset combines ground-based observations, atmospheric reanalysis, emission inventory, and other techniques to obtain nationwide seamless ground PM2.5 data from 2000 to the present. The main scope is the entire China area, the spatial resolution is 1 km, the time resolution is daily, monthly, and yearly, and the unit is µg/m3.
E.2 External experiment settings of GeoHG.
We introduce the best hyperparameter configurations for each task as below:
  • For the Carbon dataset, the hypergate $\theta_{Env}$ is set as 0.6 while $\theta_{Soc}$ is set as 0.9. The number of layers of the graph convolution block is 3, and the dimension of the hidden layer is 64.
  • For the Population dataset, the hypergate $\theta_{Env}$ is set as 0.2 while $\theta_{Soc}$ is set as 0.9. The number of layers of the graph convolution block is 3, and the dimension of the hidden layer is 64.
  • For the GDP dataset, the hypergate $\theta_{Env}$ is set as 0.4 while $\theta_{Soc}$ is set as 1.2. The number of layers of the graph convolution block is 3, and the dimension of the hidden layer is 64.
  • For the Night Light dataset, the hypergate $\theta_{Env}$ is set as 0.2 while $\theta_{Soc}$ is set as 0.9. The number of layers of the graph convolution block is 3, and the dimension of the hidden layer is 64.
  • For the $PM_{2.5}$ dataset, the hypergate $\theta_{Env}$ is set as 0.8 while $\theta_{Soc}$ is set as 0.6. The number of layers of the graph convolution block is 3, and the dimension of the hidden layer is 64.
E.3 Evaluation metrics.
We employ Mean Absolute Error (MAE), rooted mean squared error (RMSE), and coefficient of determination ($R^{2}$) as evaluation metrics, these metrics are calculated as below:
$\operatorname{MAE}(y,\hat{y})$ $=\frac{1}{|y|}\sum_{i=1}^{|y|}\left|y_{i}-\hat{y}_{i}\right|,$ (13)
$\operatorname{RMSE}(y,\hat{y})$ $=\sqrt{\frac{1}{|y|}\sum_{i=1}^{|y|}\left(y_{i}-\hat{y}_{i}\right)^{2}}$ (14)
$\operatorname{R^{2}}$ $=1-\frac{\sum_{i=1}^{n}\left(y_{i}-\hat{y}_{i}\right)^{2}}{\sum_{i=1}^{n}\left(y_{i}-\bar{y}\right)^{2}}$ (15)
Appendix F Details of Baselines
F.1 Description & Settings
. We describe the baselines as follows:
  • IDW (Inverse Distance Weighting) [Zimmerman et al., 1999] A simple and widely used spatial interpolation method that estimates values at unsampled locations by weighting the values of nearby sampled points inversely by their distance to the target location. It assumes that the influence of a point decreases monotonously with distance, making it effective when data points are dense and evenly distributed.
  • Universal Kriging [Brus and Heuvelink, 2007] A geostatistical interpolation technique that considers both the distance and the spatial correlation between data points. It incorporates a trend component into the interpolation model, allowing for more accurate predictions in the presence of spatially varying trends.
  • UrbanCLIP: [Yan et al., 2024] A model incorporating a multi-modal Large Language Model (LLM) to enhance the encoding of satellite imagery. This model generates descriptive texts for satellite images using the LLM and then fuses these texts with the images through an image-text contrastive learning-based approach to capture the complexity and diversity of geospatial areas.
  • PG-SimCLR [Xi et al., 2022]: A contrastive learning framework that introduces societal information (i.e., POI) into geospatial region representation learning from satellite imagery.
  • GeoStructural [Li et al., 2022]: The graph-based framework profiles geospatial regions by utilizing street segments as graph structure for adaptively fusing features from multi-level satellite and street-view images. For convenience, we refer to this method as GeoStructural.
  • UrbanVLP [Hao et al., 2024]: An region embedding method based on contrastive learning, which incorporates satellite imagery, street-view images, and spatial position structure. This method is further enhanced by incorporating a Large Language Model (LLM) and GeoCLIP [Vivanco Cepeda et al., 2024], resulting in improved robustness and performance.
F.2 GNNExplainer
To validate GeoHG’s efficacy in capturing higher-order geospatial relationships, we employ GNNExplainer [Ying et al., 2019], a model-agnostic technique for interpreting GNN predictions through identifying crucial subgraph structures and features. Specifically, for the carbon emission prediction task on the region #(35,65) in Guangzhou, we utilized GNNExplainer to extract the top 10 most influential edges, as shown in Table 7. Similar results for region #(43,87) are shown in Table 8. Based on Table  7, we can derive the following key findings:
Table 7: Top 10 important edges for carbon emission predictions for region #(35,65) in Guangzhou. The underlined font is the regional nodes adjacent to the target region, and the bold font refers to the distant region nodes.
Source Node Type Source Node Name Target Node Type Target Node Name Importance
region (35,65) society Food and Beverage 0.534
region (35,65) society Shopping Mall 0.511
region (35,65) environment Built-up 0.522
region (35,65) region (34,65) 0.526
region (35,65) region (36,65) 0.519
region (35,65) region (35,64) 0.522
region (35,65) region (34,64) 0.520
society Food and Beverage region (40,109) 0.495
society Food and Beverage region (47,49) 0.495
  1. 1)
    The target region exhibits strong associations with societal entity nodes like ”Food and Beverage” and ”Shopping Mall”, which are typically carbon-intensive due to factors such as energy usage (cooking, refrigeration), transportation (goods/personnel movement), and packaging consumption.
  2. 2)
    Adjacent regions’ carbon emissions emerge as highly influential features, an intuitive pattern arising from spatial proximity and potential environmental spillovers between neighboring areas.
  3. 3)
    Notably, important hyperedges also connect the target region to relatively distant areas like #(40, 109) and #(47, 49). Despite geographical separation, these regions share similar social contexts, being linked to the ”Food and Beverage” node, thereby providing relevant emission patterns to inform the prediction.
These interpretable insights validate GeoHG’s effectiveness in capturing the intricate high-order dependencies between environmental conditions, urbanization factors, economic activities, and their synergistic impact on carbon footprints across regions. By encoding such high-order interactions through graph structures, GeoHG offers a powerful inductive bias tailored for modeling complex geospatial phenomena governed by higher-order environment-society couplings.
Table 8: Top 10 important edges for carbon emission predictions for region #(43,87) in Guangzhou. The underlined font is the regional nodes adjacent to the target region, and the bold font refers to the distant region nodes.
Source Node Type Source Node Name Target Node Type Target Node Name Importance
region (43,87) society Food and Beverage 0.496
region (43,87) environment Built-up 0.495
region (43,87) environment Tree 0.519
region (43,87) region (42,87) 0.531
region (43,87) region (44,87) 0.526
region (43,87) region (43,86) 0.524
region (43,87) region (43,88) 0.520
environment Built-up region (48,54) 0.497
environment Tree region (119,7) 0.495
Appendix G Qualitative Demonstration
Figure 12: Visualization of Regions’ Similarity in Learned Embedding of Guangzhou. Darker regions indicate higher similarity with the region of interest.
Given the embedding $E_{A}$ of region A, the cosine similarity with any other region’s embedding can be calculated as follows:
$S_{C}(E_{A},E_{B})=\frac{\mathbf{E_{A}}\cdot\mathbf{E_{B}}}{\|\mathbf{E_{A}}\|\|\mathbf{E_{B}}\|}=\frac{\sum_{i=1}^{n}E_{Ai}E_{Bi}}{\sqrt{\sum_{i=1}^{n}E_{Ai}^{2}}\cdot\sqrt{\sum_{i=1}^{n}E_{Bi}^{2}}}$ (16)
where $E_{Ai}$ and $E_{Bi}$ are the $i_{th}$ components of vectors $E_{A}$ and $E_{B}$, respectively.
In Figure 12, we present the visualization results of two randomly selected regions regarding their regional similarity within the GeoHG-SSL geospatial embedding of Guangzhou. For Region (40,58), we identify it as a farmland area situated near water on an island. According to the heat map, our model has naturally discovered other regions on the same island, Jiabaosha Island, with closer regions exhibiting higher similarity. Interestingly, our model has also identified some remote islands in the north with similar environmental and societal functions, such as Dazhouwei Island and Sandy Bay. Furthermore, the western part of Guangzhou city, which is predominantly a mountainous industrial area, shows lower similarity results, aligning with this geographical characteristic.
For Region (36,63), we find that it is an urban resident area with green land near a river, named Caohe Village. From the heatmap, we can see that our model connects other residential areas exhibiting similar characteristics in Guangzhou. For example, Shilou Village and Lyv Village share similar traits, with residential settlements bordering flowing rivers. It is worth noting that, with the comprehensive representation of environmental, societal, and spatial information in the geospace and the specially designed graph structure, our models do not simplify the content region second-order region pair based solely on their intra-feature similarity. Instead, we consistently identify grouped high-order regions through the embedding results, rather than merely relying on individual high-similarity points.
Appendix H More Discussion
H.1 Limitations and Future Directions.
One limitation of our proposed method is that the complexity of the graph structure is directly proportional to the area size of the geospace, as the unit grid of the geospace is fixed to a $1km\times 1km$ square. This could present a challenge in worldwide large-scale applications, where the area of geospace may be extremely vast. In the future, we may explore the potential of modeling the geospace of the entire Earth through an adaptive grid partition and graph construction method. Furthermore, although we have emphasized and validated the efficacy and advantages of segmentation-based satellite imagery encoding, the upper limit of this approach is determined by the quality of segmentation results provided by third parties. Additionally, while we have demonstrated the effectiveness of our GeoHG in various real-world applications, further investigation is warranted to evaluate its performance in more diverse and complex scenarios.
H.2 Social Impacts.
Efficient socioeconomic indicator inference holds considerable benefits for the broader geographic and spatial computing communities. By comprehensively representing intra-region features and inter-region correlations, our proposed GeoHG framework exhibits significant potential to effect meaningful change across various domains, especially within the realms of smart cities and geoscience. The interoperability and efficiency provided by GeoHG facilitate a deeper understanding of complex geospace and the underlying mixed-order correlations inherent in the space. This comprehensive representation of regional spatial regions empowers stakeholders to monitor cities and environments more effectively, thereby making informed decisions that ultimately enhance individual and communal quality of life while fostering more resilient and sustainable environments.
Moreover, the remarkable data efficiency of GeoHG enables the community to investigate regional fine-grained climates with limited resources. For instance, in our experiment, we showcase its exceptional performance in predicting fine-grained ($1km\times 1km$) indicators for a large area ($16,401km^{2}$) using only 863 monitoring points. As human society progresses and the global environment changes, regional extreme climates, such as urban heat islands and local air pollution, continue to pose significant economic, environmental, and health challenges. Thus, effective fine-grained regional climate monitoring becomes increasingly paramount. It is noticed that the millions of lives lost annually due to local extreme heat and air pollution, we believe the enhanced geospatial region representation and data efficiency facilitated by our approach will support more effective human health protection measures and inform geopolitical decision-making, promoting improved urban and environmental management as well as extreme climate mitigation strategies.